OpenStudy (anonymous):

Which of the following equations is the result of completing the square on x2 - 6x - 9 = 0? A) (x-3)^2=0 B) (x-3)^2=9 c) (x-3)^2=18

OpenStudy (ivettef365):

(x-3)(x-3) = x^2-6x+9 and you have x^2-6x-9 = 0 so to convert x^2-6x+9 to x^2-6x-9, you need to subtract 18 so you can have x^2-6x+9-18 =0 so it looks like correct answer is C